# Order isomorphisms of sup-stable function spaces: continuous, Lipschitz, c-convex, and beyond

### v2

@@ 2.2 The $e_{x}$, the Dirac-like inf-irreducible functions of $\mathcal{G}$
Fix $x\in\mathscr{X}$ and $y\in\mathscr{Y}$, $\lambda\in\mathbb{R}$ and $\lambda^{\prime}\in\mathbb{R}$. By definition (18), we have that $e_{x}=\delta^{\top}_{x}$ and $e^{\prime}_{y}=\delta^{\top}_{y}$. Moreover $[\delta^{\top}_{x}+\lambda]$ is maximal since if there exists $h\in\mathcal{G}$ and $\alpha\in\mathbb{R}$ such that $\delta^{\top}_{x}+\lambda\leq h+\alpha$, then $h=\delta^{\top}_{x}+h(x)$, similarly $[\delta^{\top}_{y}+\lambda^{\prime}]$ is maximal. Moreover 17 guarantees that these are the only maximal Archimedean classes, and 16 shows that $J$ sends an $f\in\mathscr{F}$ with $[f]$ maximal onto a $Jf\in\mathcal{G}$ with $[Jf]$ maximal. Hence we “send Dirac masses onto Dirac masses” and 21 i) holds.
@@
We have shown in [AFG24, Proposition 6.2] that complete subspaces $\mathcal{G}\subset\mathbb{R}^{\mathscr{X}}$ that are both inf and sup-stable can always be described as the set of 1-Lipschitz functions for a weak $\overline{\mathbb{R}}$-valued metric $\tilde{d}:\mathscr{X}\times\mathscr{X}\to\overline{\mathbb{R}}$ and $\tilde{d}(z,x)=e_{x}(z)$. For inf-stable sets $\mathcal{G}$, the $e_{x}$ are (relatively)-inf-irreducible functions. It is possible sometimes to show they are the only ones even without resorting to Archimedean classes. For instance, for the set $\mathcal{G}_{\nearrow}^{\mathbb{R}}$ of increasing functions from $\overline{\mathbb{R}}$ to $\overline{\mathbb{R}}$, consider a candidate inf-irreducible $f$ and take $x_{0}\in\operatorname*{Dom}(f)\cap\mathbb{R}$, $A_{0}=\{x\in\mathbb{R}\,|\,f(x)>f(x_{0})\}$, then $f=\inf[\sup(f(x_{0}),f),f+i_{A_{0}}]$, which forces $f=e_{x^{\prime}_{0}}+\lambda_{0}$ for some $x^{\prime}_{0}\in\overline{\mathbb{R}}$, and for $x^{\prime}_{0}=+\infty$ we recover the inf-irreducible constant functions, which are Busemann points. So 21-i) holds for all (max,+)-isomorphisms $J:\mathcal{G}_{\nearrow}^{\mathbb{R}}\to\mathcal{G}_{\nearrow}^{\mathbb{R}}$ since they preserve inf-irreducibility. This is a case where there are no maximal Archimedean classes as discussed in 18. As we will see in Section 3.2 and hinted at already in 17, maximal Archimedean classes when they exist allow to separate within inf-irreducible points the $e_{x}$ from the horoboundary elements, among which the Busemann points.
@@ 2.3 Sup-irreducible functions and a Krein–Milman theorem
We have explored the inf-irreducible elements of a sup-stable $\mathcal{G}$ by studying the $e_{x}$, can we symmetrize the arguments to obtain the sup-irreducible elements? In general not, because $\mathcal{G}$ has no reason to be inf-stable. Nevertheless we can use the notion of minimal element of a subset $\mathcal{A}\subset\mathcal{G}$, i.e. the $u\in\mathcal{A}$ such that for any $v\in\mathcal{A}$ with $v\leq u$, we have $v=u$. This gives us a Krein–Milman theorem.
@@
Let $\mathcal{G}$ be a complete subspace of ¯$\overline{\mathbb{R}}^{\mathscr{X}}$ and assume that $\mathcal{G}$ is stable by decreasing nets. For all $x\in\mathscr{X}$, define $S(x):=\{u(\cdot)\in\mathcal{G}\,|\,u(x)\geq 0\}$, and let $\operatorname{Min}(S(x))$ be the set of the minimal elements of $S(x)$. Then $\operatorname{Min}(S(x))$ is non-empty and, for any $u\in\operatorname{Min}(S(x))$ and $\lambda\in\mathbb{R}$, $u(\cdot)+\lambda$ is sup-irreducible. Moreover we have that these functions sup-generate $\mathcal{G}$, i.e. $\mathcal{G}=\{\sup_{x\in\mathscr{X},u\in\operatorname{Min}(S(x))}u(\cdot)+a_{x,u}\,|\,a_{x,u}\in\mathbb{R}_{\bot}\}$.
@@
The set $\operatorname{Min}(S(x))$ is non-empty by Zorn’s lemma. Take $u\in\operatorname{Min}(S(x))$. Notice that necessarily $u(x)=0$, as otherwise $(u(\cdot)-u(x))\in S(x)$ is smaller than $u$. Fix $x\in\mathscr{X}$, $f,g\in\mathcal{G}$ such that $u=\sup(f,g)$, then w.l.o.g. $f(x)=u(x)=0$, so $f\leq u$ and $f\in S(x)$, hence $f=u$ and $u$ is sup-irreducible.
Fix $f\in\mathcal{G}$. Since $\mathcal{G}$ is stable by decreasing nets, for every $x\in\mathscr{X}$ we can select $u\in\operatorname{Min}(S(x))$ such that for all $x^{\prime}\in\mathscr{X}$, $f(x^{\prime})-f(x)\geq u(x^{\prime})$color=blue!10,size=] PCA: vérifier cette phrase. In other words, for all $x^{\prime}\in\mathscr{X}$,
@@
We say that $b:\mathscr{X}\times\mathscr{X}\to\overline{\mathbb{R}}$ is strictly tropically monotone if $b(x,x)+b(y,y)\geq b(x,y)+b(y,x)$ for all $x,y\in\mathscr{X}$, with equality iff $x=y$. This holds for $(\mathscr{X},d)$ a metric space with $b=-d$ or for $X$ a Hilbert space with $b=\langle\cdot,\cdot\rangle$.
@@
The set $Z$ is nonempty since $\mathscr{Y}\in Z$. It is also inductive since for any decreasing net $(\mathscr{Z}_{\alpha})_{\alpha\in\mathcal{A}}\in Z^{\mathscr{A}}$ we have that $\bigcap_{\alpha\in\mathcal{A}}\mathscr{Z}_{\alpha}\in Z$ and the decreasing sequence of nonempty compact sets is nonempty. Indeed by definition of $\bar{B}^{\circ}$, the function $(\bar{B}^{\circ}f=\sup_{x\in\mathscr{X}}b(x,\cdot)-f(x))$ is the supremum of continuous functions over $\mathscr{Y}$, so it is lower semicontinuous. Hence the function $(b(x,\cdot)-\bar{B}^{\circ}f)$ is upper semicontinuous. By compactness
@@
Fix for each $x$, one such $y_{\alpha,x}$. Let $\mathscr{Z}_{\cap}:=\bigcap_{\alpha\in\mathcal{A}}\mathscr{Z}_{\alpha}$. Since the net is decreasing for the inclusion, $\{y_{\alpha,x}\}_{\alpha\geq\alpha_{0}}\subset\mathscr{Z}_{\alpha_{0}}$, which is compact, so w.l.o.g. $y_{\alpha,x}\to y_{x}\in\mathscr{Z}_{\alpha_{0}}$. As this holds for any $\alpha_{0}$, by considering subnets, we obtain that $y_{x}\in\mathscr{Z}_{\cap}$. As $(b(x,\cdot)-\bar{B}^{\circ}f)$ is upper semicontinuouscolor=red!30]@SG: j’ai besoin que tu vérifies les preuves de la page 15. Très dangereux avec les subnets
@@
$f=\max\left(\sup_{y\in\bar{\mathscr{Z}}\cap\bar{U}}b(\cdot,y)-\bar{B}^{\circ}f(y),\sup_{y\in\bar{\mathscr{Z}}\cap(\mathscr{Y}\backslash U)}b(\cdot,y)-\bar{B}^{\circ}f(y)\right).$
Since $f$ is sup-irreducible, it is equal to one of the two functions, but this contradicts the minimality of $\bar{\mathscr{Z}}$. Hence $\bar{\mathscr{Z}}$ contains a single element, and $f=b(\cdot,y)-\bar{B}^{\circ}f(y)$ for $y\in\bar{\mathscr{Z}}$.color=blue!10,size=] PCA: je ne pense qu’on ait besoin de "stable by decreasing nets" jusqu’ici
Assume now that $\operatorname*{Rg}(B)$ is stable by decreasing nets, $\mathscr{X}=\mathscr{Y}$ and $b$ is strictly tropically monotone. We showed in 25 that the functions $u(\cdot)+\lambda$ were sup-irreducible, which gives one inclusion. To show the reverse inclusion, fix $x\in\mathscr{X}$, and take $u(\cdot)\in\operatorname{Min}(S(x))$ such that $b(\cdot,x)-b(x,x)\geq u(\cdot)$. Since $u\in\operatorname*{Rg}(B)$, for any $\epsilon>0$, we can find $y_{x,\epsilon}$ such that $u(x)\geq b(x,y_{x,\epsilon})-(\bar{B}^{\circ}u)(y_{x,\epsilon})\geq u(x)-\epsilon=-\epsilon$. So
@@
Let $\epsilon$ go to 00. By compactness of $\mathscr{Y}$, $y_{x,\epsilon}$ has a converging subnet converging to a $y_{x,0}$. Then, since $b(\textunderscore,\cdot)$ is continuous, evaluating the last expression at $y_{x,0}\in\mathscr{X}$, we obtain
@@
When $\mathcal{G}$ is stable by decreasing nets, the first part of 26 shows that considering a subset $\bar{\mathscr{Y}}\subset\mathscr{Y}$ such that $\{u(\cdot)+\lambda\}_{x\in X,\lambda\in\overline{\mathbb{R}},u\in\operatorname{Min}(S(x))}=\{b(\cdot,y)+\lambda\}_{y\in\bar{\mathscr{Y}},\lambda\in\overline{\mathbb{R}}}$ would give a sup-irreducible family $\{b(\cdot,y)\}_{y\in\bar{\mathscr{Y}}}$ of $\operatorname*{Rg}(B)$, which sup-generates it owing to 25. However we cannot ensure that $\bar{\mathscr{Y}}$ is compact, a problem that obviously does not happen in finite dimensions. Hence we will make sup-irreducibility an assumption.
We say that a kernel $b:\mathscr{X}\times\mathscr{Y}\to\mathbb{R}$ is fully-reduced if, for all $x\in\mathscr{X}$ and $y\in\mathscr{Y}$, $b(x,\cdot)$ and $b(\cdot,y)$ are sup-irreducible and if, for all $x_{0},x_{1}\in\mathscr{X}$, $y_{0},y_{1}\in\mathscr{Y}$ and $\lambda\in\mathbb{R}$, the implications $(b(\cdot,y_{0})=b(\cdot,y_{1})+\lambda\implies y_{0}=y_{1})$ and $(b(x_{0},\cdot)=b(x_{1},\cdot)+\lambda\implies x_{0}=x_{1})$ both hold, i.e. the rows and columns of $B$ are pairwise independent. It is sometimes possible to show this property directly as in 41 for convex lower semicontinuous functions. Otherwise the second part of 26 provides a testable inequality, satisfied for instance for any $p>0$ by $b=-d(x,y)^{p}$ over a metric space $(\mathscr{X},d)$. For $p\in(0,1]$, $b$ defines the inf-stable $\operatorname*{Rg}(B)$ of Hölder functions, which are indeed stable by decreasing nets. If $\mathscr{Y}$ is finite, $\operatorname*{Rg}(B)$ is also stable by decreasing nets.color=blue!10,size=] PCA: vérifier la dernière phrase
@@
    the two kernels are such that there exists two homeomorphisms $\tau:\mathscr{X}^{\prime}\to\mathscr{X}$ and $\sigma:\mathscr{Y}^{\prime}\to\mathscr{Y}$, and two continuous functions $\psi:\mathscr{X}^{\prime}\to\mathbb{R}$ and $\varphi:\mathscr{Y}^{\prime}\to\mathbb{R}$ satisfying
@@
Let us show the regularity of the functions, but let us avoid dealing with all functions simultaneously. We will use a translation property. Fix a $y^{\prime}_{0}\in\mathscr{Y}^{\prime}$. As in 15, $J$ allows to define a (max,+)-isomorphism $J^{\prime}$ from $\operatorname*{Rg}(B^{\prime})$ to $\operatorname*{Rg}(C^{\prime})$ where $b^{\prime}(x,\sigma(y^{\prime})):=b(x,\sigma(y^{\prime}))-b(x,\sigma(y_{0}^{\prime}))$ and $c^{\prime}(x,y^{\prime}):=c(x,y^{\prime})-c(x,y_{0}^{\prime})$. Since $J^{\prime}$ also satisfies 21-i), we can thus use (23), and obtain that $J^{\prime}f=\psi^{\prime}+f\circ\tau$ with the same $\tau$ and $\sigma$ as $J$ and some $\psi^{\prime}(\cdot)$, owing to 15. Moreover, as previously because sup-irreducibility is preserved by translations, we must have $c^{\prime}(\cdot,y^{\prime})=J^{\prime}(b^{\prime}(\cdot,\sigma(y^{\prime}))+\varphi^{\prime}(y^{\prime}))$. Hence
@@
However since the kernel $b$ is fully-reduced and $\sigma$ is bijective, we must have $x_{0}=x_{1}$. Consequently, all converging subnets have the same limit $x_{0}$. Similarly to [AB06, Lemma 2.17], if $(\tau(x^{\prime}_{\alpha}))_{\alpha\in\mathscr{A}}$ did not converge to $x_{0}$, then it would have a subnet staying away from a neighborhood of $x_{0}$, but by compactness this subnet has itself a converging subnet to $x_{0}$.color=blue!10,size=] PCA: @SG: vérifie stp We have thus shown that $\tau$ is continuous. Symmetrizing the proof, so is $\tau^{-1}$ and we get that $\tau$ and $\sigma$ are homeomorphisms. As we already established (25), since $b$ and $c$ are continuous and the bijections $\sigma$ and $\tau$ are continuous, so are $\psi$ and $\varphi$.
@@
Characterizing the (max,+)-isomorphisms provides the unexpected consequence to show that in some cases the value of “the canonical dual problem” defined below is unique:
Lemma 28 (Unique dual value).
Let $\mathcal{G}$ be a complete subspace of ¯$\overline{\mathbb{R}}^{\mathscr{X}}$. Take $f,g\in\overline{\mathbb{R}}^{\mathscr{X}}$ with $g\in\mathcal{G}$ and assume that $\mathcal{G}=\operatorname*{Rg}(B)$ and $\mathcal{G}=\operatorname*{Rg}(C)$ for two kernels $b:\mathscr{X}\times\mathscr{Y}\to\mathbb{R}$ and $c:\mathscr{X}\times\mathscr{Y}^{\prime}\to\mathbb{R}$. Assume furthermore that every (max,+)-isomorphism $J:\operatorname*{Rg}(B^{\circ})\to\operatorname*{Rg}(C^{\circ})$ is of the form (23), and there exists one such isomorphism. Consider the minimization problem $\inf_{x\in\mathscr{X}}f(x)+g(x)$, then we have
$v=\sup_{y\in\mathscr{Y}}[\inf_{z\in\mathscr{X}}[g(z)-b(z,y)]+\inf_{x\in\mathscr{X}}[f(x)+b(x,y)]]=\sup_{y^{\prime}\in\mathscr{Y}^{\prime}}[\inf_{z\in\mathscr{X}}[g(z)-c(z,y^{\prime})]+\inf_{x\in\mathscr{X}}f(x)+c(x,y^{\prime})],$ (26)
in other words, the value $v$ of the dual problem does not depend on the kernel generating $\mathcal{G}$.
Proof.
We have that $g=\bar{B}\bar{B}^{\circ}g$ as shown in [Sin97, Corollary 8.5], [AGK05, p.3], and the dual problem is obtained by permuting min and max,
$\inf_{x\in\mathscr{X}}f(x)+g(x)=\inf_{x\in\mathscr{X}}f(x)+\sup_{y\in\mathscr{Y}}b(x,y)-\bar{B}^{\circ}g(y)\geq\sup_{y\in\mathscr{Y}}-\bar{B}^{\circ}g(y)+\inf_{x\in\mathscr{X}}f(x)+b(x,y).$
By 11, we have that $\bar{C}^{\circ}=J\bar{B}^{\circ}$ where $J:\operatorname*{Rg}(B^{\circ})\to\operatorname*{Rg}(C^{\circ})$ is a (max,+)-isomorphism. By assumption, we have $Jf=\varphi+f\circ\sigma$ where $\varphi\in\mathbb{R}^{\mathscr{Y}^{\prime}}$, and the map $\sigma:\mathscr{Y}^{\prime}\rightarrow\mathscr{Y}$ is bijective. Consequently for all $h\in\overline{\mathbb{R}}^{\mathscr{X}}$
$\bar{C}^{\circ}h(y^{\prime})=J\bar{B}^{\circ}h=\varphi(y^{\prime})+\sup_{z\in\mathscr{X}}b(z,\sigma(y^{\prime}))-h(z).$
Since by 9, the kernel of the operator $\bar{C}^{\circ}$ is unique, we have that
$\bar{C}^{\circ}h=\sup_{z\in\mathscr{X}}c(z,\cdot)-h(z)\quad\text{ with }\quad c(x,y)=\varphi(y^{\prime})+b(x,\sigma(y^{\prime})).$
So the dual problem in $\bar{C}$ reads as follows, using the bijectivity of $\sigma$,
$\sup_{y^{\prime}\in\mathscr{Y}^{\prime}}\left[-\sup_{z\in\mathscr{X}}[\varphi(y^{\prime})+b(z,\sigma(y^{\prime}))-g(z)]+\inf_{x\in\mathscr{X}}[f(x)+\varphi(y^{\prime})+b(x,\sigma(y^{\prime}))]\right]\\ =\sup_{y\in\mathscr{Y}}-\bar{B}g(y)+\inf_{x\in\mathscr{X}}f(x)+b(x,y).$
Let us provide an example where dual values are different. Similarly to 22 define two kernels through the matrices
$B=\left(\begin{array}[]{ccc}0&-1&0\\ -1&0&0\\ 0&0&0\end{array}\right),\qquad C=\left(\begin{array}[]{cc}0&-1\\ -1&0\\ 0&0\end{array}\right).$
Clearly $\operatorname*{Rg}(B)=\operatorname*{Rg}(C)$ since the columns of $C$ are contained in those of $B$ and the third column of $B$ is the pointwise supremum of those of $C$. However for $f=g=(0,0,0)=C(0,0)$, we do have $g\in\operatorname*{Rg}(C)$, $0=\inf_{x}(f+g)$ and
$v_{B}=\sup_{y\in\mathscr{Y}}[-\sup_{z\in\mathscr{X}}b(z,y)+\inf_{x\in\mathscr{X}}b(x,y)]]=\max(-0-1,-0-1,0)=0,\quad v_{C}=\max(-0-1,-0-1)=-1.$

### v3

@@ 2.2 The $e_{x}$, the Dirac-like inf-irreducible functions of $\mathcal{G}$
Fix $x\in\mathscr{X}$ and $y\in\mathscr{Y}$, $\lambda\in\mathbb{R}$ and $\lambda^{\prime}\in\mathbb{R}$. By definition (18), we have that $e_{x}=\delta^{\top}_{x}$ and $e^{\prime}_{y}=\delta^{\top}_{y}$. Moreover $\delta^{\top}_{x}+\lambda]$ is maximal since if there exists $h\in\mathcal{G}$ and $\alpha\in\mathbb{R}$ such that $\delta^{\top}_{x}+\lambda\leq h+\alpha$, then $h=\delta^{\top}_{x}+h(x)$, similarly $\delta^{\top}_{y}+\lambda^{\prime}]$ is maximal. Moreover ˜17 guarantees that these are the only maximal Archimedean classes, and ˜16 shows that $J$ sends an $f\in\mathscr{F}$ with $[f]$ maximal onto a $Jf\in\mathcal{G}$ with $[Jf]$ maximal. Hence we “send Dirac masses onto Dirac masses” and ˜21 i) holds.
@@
We have shown in [AFG24, Proposition 6.2] that complete subspaces $\mathcal{G}\subset\mathbb{R}^{\mathscr{X}}$ that are both inf and sup-stable can always be described as the set of 1-Lipschitz functions for the weak $\overline{\mathbb{R}}$-valued metric $\tilde{d}:\mathscr{X}\times\mathscr{X}\to\overline{\mathbb{R}}$ with $\tilde{d}(z,x)=e_{x}(z)$. For inf-stable sets $\mathcal{G}$, the $e_{x}$ are (relatively)-inf-irreducible functions. It is possible sometimes to show they are the only ones even without resorting to Archimedean classes. For instance, for the set $\mathcal{G}_{\nearrow}^{\mathbb{R}}$ of increasing functions from $\overline{\mathbb{R}}$ to $\overline{\mathbb{R}}$, consider a candidate inf-irreducible $f$ and take $x_{0}\in\operatorname*{Dom}(f)\cap\mathbb{R}$, $A_{0}=\{x\in\mathbb{R}\,|\,f(x)>f(x_{0})\}$, then $f=\inf[\sup(f(x_{0}),f),f+\delta_{A_{0}}]$ where $\delta_{A_{0}}$ is the indicator function of $A_{0}$. This forces $f=e_{x^{\prime}_{0}}+\lambda_{0}$ for some $x^{\prime}_{0}\in\overline{\mathbb{R}}$, and for $x^{\prime}_{0}=+\infty$ we recover the inf-irreducible constant functions, which are Busemann points. So ˜21-i) holds for all (max,+)-isomorphisms $J:\mathcal{G}_{\nearrow}^{\mathbb{R}}\to\mathcal{G}_{\nearrow}^{\mathbb{R}}$ since they preserve inf-irreducibility. This is a case where there are no maximal Archimedean classes as discussed in Remark 18. As we will see in Section˜3.2 and hinted at already in ˜17, maximal Archimedean classes when they exist allow to separate within inf-irreducible points the $e_{x}$ from the horoboundary elements, among which the Busemann points.
@@ 2.3 Sup-irreducible functions and a Krein–Milman theorem
We have explored the inf-irreducible elements of a sup-stable $\mathcal{G}$ by studying the $e_{x}$, can we symmetrize the arguments to obtain the sup-irreducible elements? In general not, because $\mathcal{G}$ has no reason to be inf-stable. Nevertheless we can use the notion of minimal element of a subset $\mathcal{A}\subset\mathcal{G}$, i.e. the $u\in\mathcal{A}$ such that for any $v\in\mathcal{A}$ with $v\leq u$, we have $v=u$. We say that $\mathcal{G}$ is stable by decreasing nets if the infimum of any decreasing net of elements of $\mathcal{G}$ also belongs to $\mathcal{G}$. Considering minimal elements gives us a Krein–Milman theorem.
@@
Let $\mathcal{G}$ be a complete subspace of $\overline{\mathbb{R}}^{\mathscr{X}}$ and assume that $\mathcal{G}$ is stable by decreasing nets. For all $x\in\mathscr{X}$, define $S(x):=\{u(\cdot)\in\mathcal{G}\,|\,u(x)\geq 0\}$, and let $\operatorname{Min}(S(x))$ be the set of the minimal elements of $S(x)$. Then $\operatorname{Min}(S(x))$ is non-empty and, for any $u\in\operatorname{Min}(S(x))$ and $\lambda\in\mathbb{R}$, $u(\cdot)+\lambda$ is sup-irreducible. Moreover we have that these functions sup-generate $\mathcal{G}$, i.e. $\mathcal{G}=\{\sup_{x\in\mathscr{X},u\in\operatorname{Min}(S(x))}u(\cdot)+a_{x,u}\,|\,a_{x,u}\in\mathbb{R}_{\bot}\}$.
@@
Since $\mathcal{G}$ is stable by decreasing nets, it follows from Zorn’s lemma that $\operatorname{Min}(S(x))$ is non-empty. Take $u\in\operatorname{Min}(S(x))$. Notice that necessarily $u(x)=0$, as otherwise $(u(\cdot)-u(x))\in S(x)$ is smaller than $u$. Fix $x\in\mathscr{X}$, $f,g\in\mathcal{G}$ such that $u=\sup(f,g)$, then w.l.o.g. $f(x)=u(x)=0$, so $f\leq u$ and $f\in S(x)$, hence $f=u$ and $u$ is sup-irreducible.
Fix $f\in\mathcal{G}$. Since $\mathcal{G}$ is stable by decreasing nets, for every $x\in\mathscr{X}$ we can select $u\in\operatorname{Min}(S(x))$ such that for all $x^{\prime}\in\mathscr{X}$, $f(x^{\prime})-f(x)\geq u(x^{\prime})$. In other words, for all $x^{\prime}\in\mathscr{X}$,
@@
We say that $b:\mathscr{X}\times\mathscr{X}\to\overline{\mathbb{R}}$ is strictly tropically monotone if $b(x,x)+b(y,y)\geq b(x,y)+b(y,x)$ for all $x,y\in\mathscr{X}$, with equality iff $x=y$. This holds for $(\mathscr{X},d)$ a metric space with b=db=-d or for $X$ a Hilbert space with $b=\langle\cdot,\cdot\rangle$.
@@
The set $Z$ is nonempty since $\mathscr{Y}\in Z$. We next show that it is inductive in the sense of Bourbaki, i.e. any totally ordered subset has a maximal element (here a minimal one). Consider a decreasing net $(\mathscr{Z}_{\alpha})_{\alpha\in\mathcal{A}}\in Z^{\mathscr{A}}$. The intersection of compact sets $\bigcap_{\alpha\in\mathcal{A}}\mathscr{Z}_{\alpha}\in Z$ is compact. Moreover, by definition of $\bar{B}^{\circ}$, the function $(\bar{B}^{\circ}f=\sup_{x\in\mathscr{X}}b(x,\cdot)-f(x))$ is the supremum of continuous functions over $\mathscr{Y}$, so it is lower semicontinuous. Hence the function $(b(x,\cdot)-\bar{B}^{\circ}f)$ is upper semicontinuous. By compactness
@@
Fix for each $x$, one such $y_{\alpha,x}$. Let $\mathscr{Z}_{\cap}:=\bigcap_{\alpha\in\mathcal{A}}\mathscr{Z}_{\alpha}$. Since the net is decreasing with respect to the inclusion, $\{y_{\alpha,x}\}_{\alpha\geq\alpha_{0}}\subset\mathscr{Z}_{\alpha_{0}}$, which is compact, so w.l.o.g. $y_{\alpha,x}\to y_{x}\in\mathscr{Z}_{\alpha_{0}}$. As this holds for any $\alpha_{0}$, by considering subnets, we obtain that $y_{x}\in\mathscr{Z}_{\cap}$. As $(b(x,\cdot)-\bar{B}^{\circ}f)$ is upper semicontinuous,
@@
[…]
Since $f$ is sup-irreducible, it is equal to one of the two functions, but this contradicts the minimality of $\bar{\mathscr{Z}}$. Hence $\bar{\mathscr{Z}}$ contains a single element, and $f=b(\cdot,y)-\bar{B}^{\circ}f(y)$ for $y\in\bar{\mathscr{Z}}$.
Assume now that $\operatorname*{Rg}(B)$ is stable by decreasing nets, $\mathscr{X}=\mathscr{Y}$ and $b$ is strictly tropically monotone. We showed in ˜25 that the functions $u(\cdot)+\lambda$ were sup-irreducible, which gives one inclusion. To show the reverse inclusion, fix $x\in\mathscr{X}$, and take $u(\cdot)\in\operatorname{Min}(S(x))$ such that $b(\cdot,x)-b(x,x)\geq u(\cdot)$. Since $u\in\operatorname*{Rg}(B)$, for any $\epsilon>0$, we can find $y_{x,\epsilon}$ such that $u(x)\geq b(x,y_{x,\epsilon})-(\bar{B}^{\circ}u)(y_{x,\epsilon})\geq u(x)-\epsilon=-\epsilon$. So
@@
[…]
@@
When $\mathcal{G}$ is stable by decreasing nets, the first part of ˜26 shows that considering a subset $\bar{\mathscr{Y}}\subset\mathscr{Y}$ such that $\{u(\cdot)+\lambda\}_{x\in X,\lambda\in\overline{\mathbb{R}},u\in\operatorname{Min}(S(x))}=\{b(\cdot,y)+\lambda\}_{y\in\bar{\mathscr{Y}},\lambda\in\overline{\mathbb{R}}}$ would give a sup-irreducible family $\{b(\cdot,y)\}_{y\in\bar{\mathscr{Y}}}$ of $\operatorname*{Rg}(B)$, which sup-generates it owing to ˜25. However we cannot ensure that $\bar{\mathscr{Y}}$ is compact, a problem that obviously does not happen in finite dimensions. Hence we will make sup-irreducibility an assumption.
We say that a kernel $b:\mathscr{X}\times\mathscr{Y}\to\mathbb{R}$ is fully-reduced if, for all $x\in\mathscr{X}$ and $y\in\mathscr{Y}$, $b(x,\cdot)$ and $b(\cdot,y)$ are sup-irreducible and if, for all $x_{0},x_{1}\in\mathscr{X}$, $y_{0},y_{1}\in\mathscr{Y}$ and $\lambda\in\mathbb{R}$, the implications $(b(\cdot,y_{0})=b(\cdot,y_{1})+\lambda\implies y_{0}=y_{1})$ and $(b(x_{0},\cdot)=b(x_{1},\cdot)+\lambda\implies x_{0}=x_{1})$ both hold, i.e. the rows and columns of $B$ are pairwise independent. It is sometimes possible to show this property directly as in ˜40 for convex lower semicontinuous functions. Otherwise the second part of ˜26 provides a testable inequality, satisfied for instance for any $p>0$ by $b=-d(x,y)^{p}$ over a metric space $(\mathscr{X},d)$. For $p\in(0,1]$, $b$ defines the inf-stable $\operatorname*{Rg}(B)$ of Hölder functions, which are indeed stable by decreasing nets. If $\mathscr{Y}$ is finite, $\operatorname*{Rg}(B)$ is also stable by decreasing nets.
@@
    the two kernels are such that there exist two homeomorphisms $\tau:\mathscr{X}^{\prime}\to\mathscr{X}$ and $\sigma:\mathscr{Y}^{\prime}\to\mathscr{Y}$, and two continuous functions $\psi:\mathscr{X}^{\prime}\to\mathbb{R}$ and $\varphi:\mathscr{Y}^{\prime}\to\mathbb{R}$ satisfying
@@
Let us show the regularity of the functions, but let us avoid dealing with all functions simultaneously. We will use a translation property. Fix a $y^{\prime}_{0}\in\mathscr{Y}^{\prime}$. As in ˜15, $J$ allows to define a (max,+)-isomorphism $J^{\prime}$ from $\operatorname*{Rg}(B^{\prime})$ to $\operatorname*{Rg}(C^{\prime})$ where $b^{\prime}(x,\sigma(y^{\prime})):=b(x,\sigma(y^{\prime}))-b(x,\sigma(y_{0}^{\prime}))$ and $c^{\prime}(x,y^{\prime}):=c(x,y^{\prime})-c(x,y_{0}^{\prime})$. Since $J^{\prime}$ also satisfies ˜21-i), we can thus use (23), and obtain that $J^{\prime}f=\psi^{\prime}+f\circ\tau$ with the same $\tau$ and $\sigma$ as $J$ and some $\psi^{\prime}(\cdot)$, owing to ˜15. Moreover, as previously because sup-irreducibility is preserved by translations, we must have $c^{\prime}(\cdot,y^{\prime})=J^{\prime}(b^{\prime}(\cdot,\sigma(y^{\prime}))+\varphi^{\prime}(y^{\prime}))$. Hence
@@
However since the kernel $b$ is fully-reduced and $\sigma$ is bijective, we must have $x_{0}=x_{1}$. Consequently, all converging subnets have the same limit $x_{0}$. Similarly to [AB06, Lemma 2.17], if $(\tau(x^{\prime}_{\alpha}))_{\alpha\in\mathscr{A}}$ did not converge to $x_{0}$, then it would have a subnet staying away from a neighborhood of $x_{0}$, but by compactness this subnet has itself a converging subnet to $x_{0}$. We have thus shown that $\tau$ is continuous. Symmetrizing the proof, so is $\tau^{-1}$ and we get that $\tau$ and $\sigma$ are homeomorphisms. As we already established (25), since $b$ and $c$ are continuous and the bijections $\sigma$ and $\tau$ are continuous, so are $\psi$ and $\varphi$.